OpenStudy (anonymous):

simplify the expression 0.1y^2+0.5y^2-0.3y^2

OpenStudy (anonymous):

(0.1+0.5-0.3)y^2

OpenStudy (anonymous):

= y^2 *(0.1+0.5-0.3) = y^2 *(0.3)

OpenStudy (anonymous):

\[0.1y^2 + 0.5y^2 - 0.3y^2 = y^2(0.1+0.5-0.3) = 0.3y^2\]